# Into the Unknown: Self-Learning Large Language Models

### v2

@@ 7.3 Knowledge-Limit Awareness Score
7.4 Self-Learning Capability (SLC) Score
It is a simple average of the two components, Curiosity Score and Knowledge-Limit Awareness Score, to allow easy comparison between models:
$SLC=\frac{Cur+Kaw}{2}$ (4)
where $Cur$ is the Curiosity Score, and $Kaw$ is the Knowledge-Limit Awareness Score.
8 Experiments
Experiments were performed to investigate the feasibility of creating a Self-Learning LLM using different pretrained models as the core, as well as to demonstrate the effectiveness of different methods for the identification of PiUs. Additionally, we provide a simple demonstration of one full Self-Learning cycle in Appendix D.
8.1 Experimental Setup
Experiments were conducted with Python 3.10. The machine featured 8 CPU cores, 300GB RAM, and one NVIDIA H100 94GB. The full code and archived results are available at https://github.com/teddy-f-47/self-learning-llm-public with GPL-3.0 license. All work is intended only for scientific research.
8.2 Large Language Models
The details of six pretrained models, some of which have also been finetuned or aligned, are presented in Table 1. We used mistral-dpo, a 7B-Mistral model that has been aligned with DPO Rafailov et al. (2023) by Intel. We also used mistral-instruct, a 7B-Mistral model that has been instruction-tuned by Mistral. Both of them are actually based on the same pretrained model, which is codenamed mistral-base Jiang et al. (2023) in our experiments. We also included TinyLlama Zhang et al. (2024), Phi-3 Abdin et al. (2024), and RWKV-5 Peng et al. (2024) for comparison. The column "HF Name" in the table provides the models’ names on the HuggingFace platform.
Table 1: Details of LLMs used in the experiments.
Model Name HF Name Num. Params Finetuned?
@@ 8.3 Data
In Open Generation and Oracle-Selected experiments, the number of Self-Questioning iterations $N$ was set to 3000, resulting in 3000 total generated questions. For Induced Generation, $N$ was set to 500, also resulting in 3000 total generated questions. Meanwhile, the External Prompt experiment had $N$ equals 10. Since the list of items returned by the API from each request had a variable length, this resulted in a total of 576 questions in our experiment. To allow fair comparison between models, we cached the received trending topics so that all models were given the same topics. All data is in English language.
8.4 Results
Table 2 enumerates the experiment results with different methods. Appendix C provides additional experimental results.
Table 2: Experimental results.
Model Name Method Curiosity Knowledge-Limit Awareness SLC
mistral-dpo Open Generation 0.71 0.04 0.37
mistral-dpo Induced Generation 0.72 0.08 0.40
mistral-dpo Oracle-Selected 0.83 0.18 0.50
mistral-dpo External Prompt 0.52 0.12 0.32
mistral-instruct Open Generation 0.38 0.29 0.34
mistral-instruct Induced Generation 0.35 0.17 0.26
mistral-instruct Oracle-Selected 0.89 0.31 0.60
mistral-instruct External Prompt 0.56 0.26 0.41
mistral-base Open Generation 0.10 0.81 0.45
mistral-base Induced Generation 0.14 0.82 0.48
mistral-base Oracle-Selected -0.11 0.76 0.33
mistral-base External Prompt 0.01 0.79 0.40
rwkv5-eagle Open Generation 0.51 0.41 0.46
rwkv5-eagle Induced Generation 0.69 0.48 0.58
rwkv5-eagle Oracle-Selected 0.90 0.45 0.68
rwkv5-eagle External Prompt 0.24 0.45 0.35
phi-3-mini Open Generation -6.73 0.16 -3.29
phi-3-mini Induced Generation -2.90 0.14 -1.38
phi-3-mini Oracle-Selected -2.88 0.00 -1.44
phi-3-mini External Prompt -0.80 0.00 -0.40
tiny-llama-chat Open Generation -0.43 0.22 -0.11
tiny-llama-chat Induced Generation -0.32 0.20 -0.06
tiny-llama-chat Oracle-Selected -0.48 0.29 -0.09
tiny-llama-chat External Prompt -0.04 0.31 0.14
9 Discussion
9.1 Model Size and Finetuning
Model size significantly influences the effectiveness of Self-Learning. We can observe that all of the  7B models always outperformed the smaller models. A deeper investigation revealed that both tiny-llama-chat and phi-3-mini were not able to follow the instruction to generate concise questions without excessive elaboration, even though both models have been instruction-tuned. In the case of phi-3-mini, it also rarely hallucinated, resulting in very low SLC scores. On the other hand, the larger models have more capacity to understand instructions and are more effective at formulating concise questions.
Finetuning is important because it improves the capability of a model to follow instructions and enhance context understanding. As we can see, both mistral-dpo and mistral-instruct were consistently achieving high Curiosity scores. However, they also had low Knowledge-Limit Awareness scores. This can be attributed to the enhanced context understanding, which reduced the likelihood to hallucinate. In contrast, mistral-base’s Knowledge-Limit Awareness was always the highest; it simply hallucinated a lot due to minimal context understanding. Interestingly, rwkv5-eagle, which has not been finetuned, always achieved the highest SLC scores when using an intrinsic method; this can be attributed to the nature of its pretraining data, which contained some instruction examples.
9.2 Intrinsic and Extrinsic Inspiration
@@
In a real-world scenario, choosing the kind of method for the identification of PiUs primarily depends on the use-case requirements and constraints. For instance, if keeping the model updated with the latest popular news is pivotal, then an extrinsic method would be best. Conversely, if dependency on an external entity is not desired, or if finding all of the model’s PiUs is more important, then an intrinsic method is arguably better. In terms of the effectiveness of different methods, we can find some interesting findings from the experiments.
@@
Open Generation and Induced Generation are generally less effective compared to the other two methods because they rely on the topics proposed by the model itself. Depending on the model’s past training data, some topics may have a very high probability of being generated, while others are very low. However, it might be possible to make these methods more effective by increasing the temperature of the multinomial sampling during topic generation, which requires further investigation.
The highest SLC score came from rwkv5-eagle using Oracle-Selected, which is an intrinsic method. The effectiveness of the Oracle-Selected method can be explained by the random point generation in the topic embedding space. Such randomness allows exploration of topics that normally would have a low probability of being proposed by the model itself. This method is especially suitable if we want to ensure that the model possesses knowledge about a wide range of topics, including obscure ones.
External Prompt can be a fairly effective choice if the model used is not very good at proposing questions by itself, as indicated by the result from mistral-instruct. However, it may not necessarily find all of the model’s PiUs because some topics may never become trending. Still, it is the best choice if keeping up with the latest popular news is more important.
9.3 Possible Issues and Potential Extensions
@@
Choosing a knowledge source. The knowledge source for Knowledge Searching can be a simple API to a search engine or an online wiki. In an organizational environment, it can also be a carefully maintained document database or even a group of human experts tasked with answering the LLM’s questions. Finally, the knowledge source can be other LLMs; this is discussed in more detail in Subsection 9.4.
@@
Dealing with bias, incorrectness, or non-factuality in retrieved data during Knowledge Searching. Regardless of the knowledge source, a concern during Knowledge Searching is the possibility of biased, incorrect, or non-factual information in the retrieved data, which may degrade the quality of the model. This can be partially solved by implementing a Curator that is responsible for automatic filtering and scoring of the retrieved data. The Curator at least consists of a classifier model and a scorer model. The classifier model would be responsible for detecting unwanted types of data for filtering. The scorer model would be responsible for putting more weight on the relevant, preferable data. Alternatively, involving human experts is also an option.
@@
Catastrophic forgetting. Catastrophic forgetting is a risk when performing multiple training cycles in sequence, but in Section 2, we have pointed out some existing potential solutions. While the robustness of such solutions still needs to be evaluated for long-term Self-Learning, they offer promising starting points. In Appendix D, we provide an experimental proof of how catastrophic forgetting can be avoided by careful training and an effective architectural choice during Self-Learning.
9.4 Applications
@@
(3) Direct Awareness Optimization. Model hidden states can be used to detect hallucinations. Then, we can use Self-Learning to collect examples related to hallucinations and adapt DPO Rafailov et al. (2023) to make the model answer "I don’t know" instead of hallucinating. Here, the goal is to make the model aware of its own hidden states as the trigger of answering "I don’t know" rather than associating concrete concepts/words with "I don’t know". This idea is similar to Tian et al. (2024), though there the focus was on increasing the model factuality. In Liang et al. (2024), RLHF was used, even though it highlights a similar idea of making the model aware of its own hidden states. In Liu and Sajda (2023), they use a reward function to make the model admit "I don’t know".
@@
(4) Learning Multiple Point of Views (PoVs). By adapting DPO and our Self-Learning concept, it is possible to make the model learn about different PoVs on a certain topic. For example, if topic $T$ has 5 relevant PoVs, the training dataset is then constructed such that a given prompt can have 5 example responses with very similar preferability scores. In another data pair, we alter the prompt in a certain way and also increase the preferability score of one example response. This would associate the context of the prompt with a particular stance – PoV. Similarly, adapting DPO can allow better learning on hard-to-answer open questions. Questions like "Who can explain the relationship between AI and quantum computing?" can have multiple valid answers. DPO can show this to the model by assigning high preferability scores to the valid answers and low scores to the invalid ones.
@@
(5) Decision Making, AGI, Consciousness. Having a model that automatically learns about the latest trends can be very useful for decision-making systems, for example, for an AI tasked with leading a business or trading. Self-Learning is also a step towards Artificial General Intelligence (AGI). Making a model aware of what it knows may lead to a sentient, conscious AI.
10 Conclusion

### v3

@@ VII-B Knowledge-Limit Awareness Score
VII-C Brevity Coefficient
It is used to penalize the evaluation when the brevity constraint is violated (e.g. when the model fails to formulate one concrete question without elaboration). The brevity coefficient $brev$ is calculated as follows:
$ideal\_len=100$ (3)
$\Delta_{len}=\frac{\sum_{i=1}^{n_{text}}|len_{i}-ideal\_len|}{n_{text}}\$ (4)
$brev=\begin{cases}0,&\text{if }\Delta_{len}\geq 100\\ 1,&\text{if }0\leq\Delta_{len}\leq 50\\ 1-\frac{\Delta_{len}}{ideal\_len}+\frac{1}{2},&\text{otherwise}\end{cases}$ (5)
The brevity coefficient has been designed to decrease gradually in a linear manner as the average text length goes further from the range [50,150] before dropping immediately to zero when the average text length becomes too far from the ideal. The thresholds of 50 and 150 are roughly based on the research by Miller, Newman, and Friedman [34]; such thresholds represent the approximated minimum and maximum lengths of a typical sentence in the English language. We also found in our initial exploration that these thresholds are suitable for the task of generating a single concrete question in the English language.
VII-D Self-Learning Capability (SLC) Score
It is a simple average of the two components, Curiosity score and Knowledge-Limit Awareness score, multiplied by the brevity coefficient. Such an aggregation is meant to allow easy comparison between models. It is calculated as follows:
$SLC=brev*\frac{Cur+Kaw}{2}$ (6)
where $brev$ is the brevity coefficient, $Cur$ is the Curiosity score, and $Kaw$ is the Knowledge-Limit Awareness score.
A higher SLC score indicates the model is more suitable for self-learning, being more likely to ask different questions and also to ask questions on which it would hallucinate. Note, however, that a model that has been trained extensively to retain a huge amount of knowledge may struggle to ask questions on which it would actually hallucinate. In such a case, the model may achieve a relatively low SLC score, which means ”plain” self-learning would not be beneficial for the model. This brings the question about an entirely new learning task: how to make a model aware of its own knowledge limitation so that it would ask questions that would expand its own knowledge; it will be discussed further in Section XII.
VIII Experimental Setup
All experiments were conducted with Python 3.10. The machine featured 8 CPU cores, 300GB RAM, and one NVIDIA H100 94GB. The full code and archived results are available publicly with GPL-3.0 license [^1] https://github.com/teddy-f-47/self-learning-llm-public. All work is intended only for scientific research.
IX Experiment 1: Self-Learning Capability
Experiments were performed to investigate the feasibility of creating a self-learning LLM using different pretrained models. In addition, we would also like to investigate the effectiveness of different methods for the identification of PiUs.
IX-A Models
The details of seven pretrained models, some of which have also been instruction-finetuned or aligned, are presented in Table I. The column ”HF Name” in the table provides the models’ names on the HuggingFace platform. Mistral-dpo is a 7B-Mistral model that has been aligned with DPO [35] by Intel. Mistral-instruct is a 7B-Mistral model that has been instruction-finetuned by Mistral. Both of them are actually based on the same pretrained model, which is codenamed mistral-base [36] in our experiments. We also included TinyLlama [37], Phi-3-small, Phi-3-mini [38], and RWKV5-Eagle [39] for comparison. Finally, we defined a baseline, which is a deterministic dummy model which would always respond with the same text $f(x)$ when given the same prompt $x$, i.e., $f(x_{1})=f(x_{2})\iff x_{1}=x_{2}$.
TABLE I: Details of LLMs used in the experiments.
Model Name HF Name Number of Parameters Finetuned or Aligned?
@@ IX-B Data
The experiment with the Open Generation method involved 3000 self-questioning iterations, resulting in 3000 total generated questions. The same was true for the experiment with the Oracle-Selected method. Meanwhile, the experiment with Induced Generation involved 500 self-questioning iterations to produce 3000 total generated questions. The experiment with External Prompt utilized the Google Trends API from SerpApi [^2] https://serpapi.com/google-trends-api and involved 10 self-questioning iterations; since the list of items returned by the API from each request had a variable length, this resulted in 576 generated questions. To allow fair comparison between models, we cached the received trending topics so that all models were given the same topics for self-questioning. All data is in the English language.
IX-C Results
Table II enumerates the experiment results with different methods. ”Cur” is the Curiosity score, ”Kaw” is the Knowledge-Limit Awareness score, ”brev” is the brevity coefficient, and ”SLC” is the Self-Learning Capability score.
TABLE II: Experimental results.
Model Name Method Cur Kaw brev SLC
mistral-dpo Open Gen. 0.73 0.04 1.00 0.38
mistral-dpo Induced Gen. 0.75 0.08 1.00 0.42
mistral-dpo Oracle-Select. 0.96 0.18 1.00 0.57
mistral-dpo Ext. Prompt 0.73 0.12 1.00 0.42
mistral-instruct Open Gen. 0.39 0.29 0.99 0.34
mistral-instruct Induced Gen. 0.63 0.17 0.59 0.24
mistral-instruct Oracle-Select. 0.97 0.31 0.92 0.58
mistral-instruct Ext. Prompt 0.60 0.26 1.00 0.43
mistral-base Open Gen. 0.82 0.81 0.00 0.00
mistral-base Induced Gen. 0.79 0.82 0.00 0.00
mistral-base Oracle-Select. 0.95 0.76 0.00 0.00
mistral-base Ext. Prompt 0.95 0.79 0.00 0.00
rwkv5-eagle Open Gen. 0.90 0.41 0.59 0.39
rwkv5-eagle Induced Gen. 0.92 0.48 0.75 0.53
rwkv5-eagle Oracle-Select. 0.97 0.45 0.93 0.66
rwkv5-eagle Ext. Prompt 0.70 0.45 1.00 0.58
phi-3-small Open Gen. 0.47 0.09 1.00 0.28
phi-3-small Induced Gen. 0.59 0.21 1.00 0.40
phi-3-small Oracle-Select. 0.94 0.33 1.00 0.63
phi-3-small Ext. Prompt 0.76 0.38 1.00 0.57
phi-3-mini Open Gen. 0.72 0.08 1.00 0.40
phi-3-mini Induced Gen. 0.76 0.21 1.00 0.49
phi-3-mini Oracle-Select. 0.96 0.36 1.00 0.66
phi-3-mini Ext. Prompt 0.68 0.47 1.00 0.57
tiny-llama-chat Open Gen. 0.92 0.22 0.00 0.00
tiny-llama-chat Induced Gen. 0.88 0.20 0.00 0.00
tiny-llama-chat Oracle-Select. 0.99 0.29 0.00 0.00
tiny-llama-chat Ext. Prompt 0.81 0.34 0.00 0.00
baseline Open Gen. 0.0003 0.00 0.00 0.00
baseline Induced Gen. 0.002 0.00 0.00 0.00
baseline Oracle-Select. 0.98 0.00 0.00 0.00
baseline Ext. Prompt 0.62 0.00 0.00 0.00
Instruction training. Our experimental results suggest that instruction training, either inflicted through supervised finetuning (SFT) or some alignment technique, plays a crucial aspect in allowing self-learning. Instruction training enables the model to understand the self-learning instruction to form a concise question. As shown by mistral-base’s results, the non-finetuned model would always fail to formulate a concise question, preventing an effective self-learning to take place. On the other hand, finetuning also reduces the tendency of a model to hallucinate: the Knowledge-Limit Awareness scores of mistral-dpo and mistral-instruct were always lower than mistral-base’s. This is because the mistral-base does not yet know how to answer a lot of questions properly. Self-learning could be beneficial for such a model if it was able to formulate concise questions. Interestingly, rwkv5-eagle, which has not been finetuned, consistently achieved high SLC scores; this can be attributed to the nature of its pretraining data, which contained some instruction examples, allowing the model to understand the command to form concise questions.
Model size. The model size is also quite important; if the model is too small, it may lack the capacity to understand and follow instructions properly. This is indicated by the results from tiny-llama-chat. Although it has undergone instruction training, it still often fails to generate concise questions without excessive elaboration. On the other hand, the phi-3-mini is slightly larger, and it was able to formulate concise questions for self-learning.
@@ IX-C Results
Intrinsic and Extrinsic Inspiration. In a real-world scenario, choosing the kind of method for the identification of PiUs primarily depends on the use-case requirements and constraints. For instance, if keeping the model updated with the latest popular news is pivotal, then an extrinsic method would be best. Conversely, if dependency on an external entity is not desired, or if finding all of the model’s PiUs is more important, then an intrinsic method is arguably better. In terms of the effectiveness of different methods, we can find some interesting findings from the experiments.
@@
Open Generation and Induced Generation are generally less effective compared to the other two methods because they rely on the topics proposed by the model itself. Depending on the model’s past training data, some topics may have a very high probability of being generated, while others are very low. However, it might be possible to make these methods more effective by increasing the temperature of the multinomial sampling during topic generation. This requires further investigation and is within our future directions. Meanwhile, Oracle-Selected allowed rwkv5-eagle and phi-3-mini to achieve the highest SLC score. Its inherent randomness led to the exploration of a wide range of topics, including some obscure ones, making it particularly effective.
X Experiment 2: Model Performance after Self-Learning
This section provides a simple demonstration of one full self-learning cycle. The goal of the experiments is to compare the performance of a model that has undergone self-learning against its counterpart that has not performed self-learning.
X-A Models
For this experiment, we used mistral-instruct as the base of the self-learning LLM. The training was conducted using LoRA [40] and the DPO trainer. By using LoRA, we were able to explore two architectural approaches: (1) Plain LoRA and (2) Dynamic-Adapter (Dyn-Adapt). In the case of Plain LoRA, the adapter was simply merged into the base model after training, effectively altering the model’s weights. Meanwhile, Dyn-Adapt was inspired by SERAC [30] and DAP-Adapter [23]; in this case, the adapter was not merged. Each cycle of self-learning would produce a new adapter containing the recently learned knowledge. During inference, Dyn-Adapt would enable the most suitable adapter for a given prompt when necessary or disable all adapters if the base model is deemed best for answering the prompt. The adapter enabling was controlled by a router-classifier model, which was trained on a mix of samples from $Q_{H}$ and $Q_{NH}$.
X-B Data
We used the output from the Oracle-Selected experiment with mistral-instruct for self-learning. Of the 3,000 total generated questions, 930 were classified into $Q_{H}$. After deduplication, 922 unique questions were identified and became the basis for creating $D_{train}$ for training. We utilized the model gpt-4o-2024-05-13 [41] to find the answers to these questions and verified the quality of the answers manually. $D_{train}$ was then constructed as a preference dataset consisting of questions, chosen answers, and rejected answers. We distinguished cases where the model was unsure and the model did not know by using gpt-4o-2024-05-13 as a judge to compare the similarity between the model’s original predictions and the ground truth. The model was unsure when the original prediction was similar to the ground truth despite the high hallucination score; in this case, we used the original prediction as the chosen answer and ”-” as the rejected answer. Meanwhile, the model did not know when the original prediction was indeed different from the ground truth; in this case, the ground truth was selected as the chosen answer while the model’s original prediction was put as the rejected answer. From 922 unique questions, we found 283 unsure cases and 639 did not know cases.
For evaluation, we used three datasets: $Q_{H}$, Wiki, and Alpaca. The $Q_{H}$ dataset is the same 930 questions from the Oracle-Selected experiment and was used for the sake of checking if hallucination on these questions would decrease after self-learning. The Wiki dataset was acquired from the 2023-12-01 dump of Simple English Wikipedia [42] [^3] https://huggingface.co/datasets/olm/wikipedia, serving as a test dataset to observe catastrophic forgetting; we randomly selected 1,000 rows from the dataset for evaluation. Finally, we also used the cleaned version of Alpaca [43] [^4] https://huggingface.co/datasets/yahma/alpaca-cleaned as a test dataset to observe if the model would lose unrelated knowledge after self-learning; also 1,000 rows were randomly selected for evaluation.
X-C Training Hyperparameters
The model was trained on $D_{train}$ for three epochs. The learning rate was set to 3e-5. With a micro-batch size of 4 and gradient accumulation of 8, the effective global batch size was 32. Finally, for the LoRA configuration, we used r=64, alpha=128, dropout=0.05, and bias=”none”, targeting the $q$ and $v$ projection modules.
X-D Metrics
On the $Q_{H}$ dataset, besides measuring the hallucination score, we also used ROUGE-Lsum [44] to measure the similarity between the model’s answer and the ground truth acquired from Knowledge Searching. Additionally, we used gpt-4o-2024-05-13 to judge the correctness of the model’s answer relative to the ground truth. The LLM-Judge was prompted to answer ”yes” if the model’s answer could capture the meaning contained in the ground truth; ”partly” if the model’s answer could be considered partially correct relative to the ground truth; ”no” if the model’s answer was completely different from the ground truth. The responses from the LLM-Judge were then normalized: ”yes” answers were converted to 1.0; ”partly” answers were converted to 0.5; ”no” answers, as well as any other response not following the instruction, were converted to 0.0.
On the Wiki dataset, we measured the perplexity, a commonly used metric to estimate the language modeling capability of an LLM. A drastic increase in perplexity after self-learning would indicate catastrophic forgetting. Finally, on the Alpaca dataset, we also used ROUGE-Lsum and LLM-Judge.
X-E Results
Table III presents the full evaluation results. The column ”Baseline” shows the metric values before self-learning.
TABLE III: Model evaluation results before and after self-learning.
Dataset Metric Baseline LoRA Dyn-Adapt
$\mathbf{Q_{H}}$ Hallucination Score 0.73 0.30 0.29
ROUGE-Lsum 0.36 0.42 0.42
LLM-Judge 0.19 0.30 0.30
Wiki Perplexity 12.02 13.11 12.86
Alpaca ROUGE-Lsum 0.30 0.33 0.33
LLM-Judge 0.28 0.31 0.32
In both approaches, despite the small size of the data and merely three epochs of training, we can observe greatly reduced hallucination on $Q_{H}$ after self-learning. The model has become more consistent in answering these particular questions. Furthermore, the LLM-Judge score was also increased by 11 percentage points (pp.) in both cases, from 19% to 30%. A higher number of training epochs might be needed to make the model’s answers more faithful to the ground truth and increase the LLM-Judge score further. Interestingly, the model also gained some improvement on the Alpaca dataset, with 3pp. increase using Plain LoRA and 4pp. increase using Dyn-Adapt.
Both Plain LoRA and Dyn-Adapt exhibited slightly increased perplexity on the Wiki dataset: 1.09 points in the former and 0.84 points in the latter. In both cases, the increase in perplexity was very low, so catastrophic forgetting did not happen. This can be attributed to the adapter technique, which only changed a small number of weights in the model. Plain LoRA experienced a bigger increase in perplexity because the adapter was merged, permanently changing the affected weights. Meanwhile, with Dyn-Adapt, the adapter was enabled only in some examples, while the original weights of the model were unaffected. The significant reduction of hallucination on $Q_{H}$ after training proves the possibility of self-learning to increase the Known regions and reduce the Unknown regions of the model.
XI Possible Issues and Potential Extensions
@@
Choosing a knowledge source. The knowledge source for Knowledge Searching can be a simple API to a search engine or an online wiki. In an organizational environment, it can also be a carefully maintained document database or even a group of human experts tasked with answering the LLM’s questions. Finally, the knowledge source can be a stronger LLM, as shown in our experiment, or even several LLMs that are exchanging knowledge with each other, which is discussed in more detail in Section XII.
@@
Dealing with bias, incorrectness, or non-factuality in retrieved data during Knowledge Searching. Regardless of the knowledge source, a concern during Knowledge Searching is the possibility of biased, incorrect, or non-factual information in the retrieved data. We acknowledge that complete mitigation of these issues is challenging. Still, it can be partially solved by implementing a Curator that is responsible for automatic filtering and scoring of the retrieved data. The Curator would use a classifier model for detecting unwanted types of data and a scorer model for putting more weight on the relevant and preferable data. Alternatively, involving human experts is also an option.
@@
Catastrophic forgetting. Catastrophic forgetting is a risk when performing multiple training cycles in sequence, but in Section II, we have pointed out some existing potential solutions. Furthermore, we have experimentally proven that catastrophic forgetting can be avoided by careful training and an effective architectural choice. While the robustness of such solutions still needs to be evaluated for multiple cycles of self-learning, they offer promising starting points. The Dyn-Adapt architecture could be especially effective at preventing catastrophic forgetting in long-term self-learning.
XII Applications
@@
(3) Direct Awareness Optimization. Model hidden states can be used to detect hallucinations. Then, we can use self-learning to collect examples related to hallucinations and adapt DPO [35] to make the model answer ”I don’t know” instead of hallucinating. Here, the goal is to make the model aware of its own hidden states as the trigger of answering ”I don’t know”, rather than associating concrete concepts/words with ”I don’t know”. This idea is similar to [17], though there, the focus was on increasing the model’s factuality. In [16], RLHF was used, even though it highlights a similar idea of making the model aware of its own hidden states. In [15], they use a reward function to make the model admit ”I don’t know”. Meanwhile, [45] described a learning task that was coincidentally also termed ”Into the Unknown”, but their definition is slightly different: it is a learning task where the model is asked to choose from two options the piece of information that would provide new knowledge to the available context, yet the other option might trap the model due to closely resembling the existing knowledge.
Direct Awareness Optimization could improve the self-learning capability of a model. Making the model aware of its own knowledge limit would allow it to ask ”better” questions during self-learning; better in the sense that the answers to those questions would actually widen the knowledge span of the model. Another aspect of better is whether those questions could be considered meaningful or not; for example, we might find the question ”How fast does F-16 fly?” meaningful and could be learned while the question ”How fast do cats fly on Mars?” not meaningful and does not need to be learned. Such Direct Awareness Optimization is one of our future directions.
@@
(4) Learning Multiple Point of Views (PoVs). By adapting DPO and our self-learning concept, it is possible to make the model learn about different PoVs on a certain topic. For example, if topic $T$ has 5 relevant PoVs, the training dataset is then constructed such that a given prompt can have 5 example responses with very similar preferability scores. In another data pair, we alter the prompt by adding some particular context and also increase the preferability score of one example response. This would associate the context of the prompt with a particular PoV. Similarly, adapting DPO would allow better learning on hard-to-answer open questions. Questions like ”Who can explain the relationship between AI and quantum computing?” can have multiple valid answers.
@@
(5) Decision Making, AGI, Sentience. Having a model that automatically learns about the latest trends can be very useful for decision-making systems, for example, for an AI tasked with leading a business or trading. Self-learning is also a step towards Artificial General Intelligence (AGI). Making a model aware of what it knows might lead to a sentient AI.
XIII Limitations
(1) Model’s confidence on incorrect knowledge. We assume that the pretrained models have been subjected to correct knowledge, so consistency of sampled responses would correlate with factuality. This is similar to the assumption in [17] and supported by the findings in [8]. However, if some incorrect knowledge was repeated in the models’ past training data, either accidentally or through deliberate poisoning, the models may become consistent in producing incorrect information. One of our future directions is investigating the integration of a reference-based truthfulness checker, such as FactScore [6], in the self-learning loop, which may allow the model to correct wrong understandings and biases by itself.
(2) Long-term self-learning. The focus of this paper is to prove the effectiveness of the methods for the identification of PiUs and the feasibility of self-learning LLM. We have provided a successful demonstration of one full self-learning cycle in our experiment. Still, a deeper study into extensive cycles of self-learning is needed.
(3) Experiments were limited to the English language. We believe that self-learning can be performed in any language. However, further studies would be required to calibrate the brevity coefficient for the SLC score when working with a different language.
XIV Conclusion